# Diffusion MRI with Machine Learning

### v2

@@ Abstract
Diffusion-weighted magnetic resonance imaging (dMRI) offers unique capabilities including noninvasive probing of brain’s tissue microstructure and structural connectivity. It is widely used for clinical assessment of brain pathologies and for neuroscience research. Analyzing the dMRI data to extract useful information for medical and scientific purposes can be challenging. The dMRI measurements often suffer from strong noise and artifacts, there is usually high inter-session and inter-scanner variability in the data, and considerable inter-subject heterogeneity in brain structure. Moreover, the relationship between measurements and the phenomena of interest can be highly complex. Recent years have witnessed increasing use of machine learning methods for dMRI analysis. This manuscript aims to assess these efforts, with a focus on methods that have addressed data preprocessing and harmonization, microstructure mapping, tractography, and white matter tract analysis. We study the main findings, strengths, and weaknesses of the existing methods and suggest topics for future research. We find that machine learning may be exceptionally suited to tackle some of the difficult tasks in dMRI analysis. However, for this to happen, several shortcomings of existing methods and critical unresolved issues need to be addressed. These include deficient evaluation practices, lack of rich training datasets and validation benchmarks, as well as model generalizability, reliability, and explainability concerns.
@@ I-A Background and motivation
Classical dMRI analysis methods have been mostly based on conventional signal processing, biophysical modeling, and numerical optimization techniques. Meanwhile, many studies have advocated for machine learning and data-driven techniques [16, 17, 18]. These techniques have become more popular in recent years. This trend has been driven by more powerful machine learning models (mostly based on deep neural networks) and a greater appreciation of the power and flexibility of these methods. Overall, these methods have been shown to possess the potential to improve the speed, accuracy, and reproducibility of various computations in dMRI analysis such as data pre-processing [19] and harmonization [20], tissue microstructure mapping [16, 21, 22], tractography [18], tract-specific analysis [17], and population studies [23]. It appears that recent studies herald a new generation of dMRI analysis methods that may significantly complement, if not at least in some cases entirely replace, the more conventional techniques. Therefore, it is good time to review, summarize, and critically assess the achievements of these works, highlight their shortcomings and limitations, and point out future work that may contribute to this new field in dMRI.
@@ I-B Scope and organization of this manuscript
The search to find the relevant works to be included in this paper was conducted in PubMed and Google Scholar. The initial search was performed in June 2023 and then repeated in June 2024. The search terms included “diffusion MRI” and “machine learning” or “deep learning”. More detailed searches were performed in reputable journals (e.g., Medical Image Analysis and IEEE Transactions on Medical Imaging) and conferences (e.g., MICCAI). Given the wide scope of this paper, we excluded the works that had little methodological novelty or limited experimental validation. Additionally, where two or more papers presented similar ideas and methods, we chose the paper with more comprehensive experimental results.
This monograph starts with a brief description of challenges in dMRI data analysis (Section II) and general overview of the reasons why machine learning may be effective in addressing these challenges (Section III). Section IV reviews prior works that have used machine learning in dMRI. The main applications considered include data preprocessing and quality enhancement, data harmonization, estimation of tissue microstructure and fiber orientation, tractography, tract-specific analysis, registration and segmentation. Figure 1 shows the outline of that section. Following that, Section V discusses the main technical considerations, challenges, and open questions. Some of the main topics discussed in that section include validation approaches, inherent limitations of machine learning, data, ground truth, and modeling considerations, as well as model explainability, uncertainty, and reliability concerns. Figure 2 shows the outline of that section. A short Conclusion section will present the closing remarks.
@@ II The challenging nature of dMRI analysis
The challenges and pitfalls of analyzing the dMRI data have been discussed in dedicated publications [9, 15, 24]. A primary source of difficulty in dMRI analysis is measurement noise and artifacts [9, 25]. Measurement noise can significantly alter the analysis results even in relatively simple computations such as diffusion tensor estimation [26, 27]. It can be difficult to suppress the noise and artifacts due to their complex distributions. The echo planar imaging method that is used to acquire dMRI measurements can give rise to eddy-current- and susceptibility-induced distortions. Subject motion during data acquisition is another persistent challenge in MRI. These factors can significantly impact the accuracy of quantitative tissue microstructure mapping and structural connectivity analysis [28, 29, 30]. Furthermore, the relationship between the dMRI measurements and the underlying microstructure of neuronal tissue is complex [31, 32]. Another source of difficulty is the measurement requirements of advanced dMRI models. Estimation of multi-compartment models of tissue microstructure and of complex white matter fiber configurations requires densely-sampled q-space data. Obtaining such measurements would require long scan times that may not be affordable in many settings. Even with a reasonably large number of measurements, standard computational methods can produce erroneous results [33]. As only one example, conventional linear and non-linear least-squares techniques for diffusion kurtosis imaging (DKI [34]) tend to produce physically implausible results [35, 36]. Streamline tractography and tract-specific analysis face inherent challenges and ambiguities such as fiber crossings and bottlenecks [24, 37, 38, 39], which can give rise to high false positive rates and erroneous results. Cross-subject comparisons and population studies can also be significantly hampered by high inter-scanner data heterogeneity [40, 41]. Furthermore, most applications are also challenged by paucity of ground truth and lack of universally accepted performance metrics.
@@ IV-F Tractography and structural connectivity
Machine learning may offer a framework for developing better tractography algorithms. They are model-free and allow for a seamless integration of anatomical priors and other sources of information that may be useful for tractography. Some of these information, such as the direction of previous tractography steps, have been used in conventional tractography methods in the past [227, 228]. Nonetheless, machine learning methods offer higher flexibility in integrating these various inputs into a unified model and to optimize the model to reduce the tractography error with respect to all inputs jointly. As a result, they have the potential to have lower false positive rates, to be more robust to measurement noise, and to overcome the inherent biases of classical tractography algorithms [18, 229, 230]. Although various methods such as self-organizing maps [231] and random forests [229] have been successfully used for tractography, machine learning-based tractography techniques have increasingly relied on neural networks [18, 229, 232, 233]. Table VI lists some of the works that have developed machine learning methods for streamline tractography. We summarize the main technical aspects of these methods and their reported results below.
@@
Jörgens et al. [236] MLP (D) Raw diffusion signal & direction of the two preceding streamline steps. This work reports extensive experiments to determine the optimal input and network architectures.
Wegmayr et al. [59] MLP (D) Raw diffusion signal in a $3^{3}$-voxel neighborhood & direction of the two preceding streamline steps. Better results than standard tractography methods when trained on small datasets. The method is robust to noisy training data.
Cai et al. [61] Convolutional-recurrent neural network Anatomical MRI and streamline memory. The variability in tractography and structural connectivity measures computed based on T1-weighted MRI is similar to that for standard dMRI-based results.
Track-to-Learn [237] Deep reinforcement learning (D & P) FOD at the current voxel and its six neighbors & direction of the past four streamline steps. A reinforcement learning approach can be competitive with supervise learning methods while also being more generalizable to unseen or out-of-distribution data.
@@
Entrack [238] MLP (P) Direction of the three last streamline steps & FOD, represented as SH of order 4, in a voxel neighborhood of size 3 voxels. The method achieves competitive results compared with a range of conventional and machine learning methods on independent synthetic test data.
@@ IV-F3 Evaluations and main results
Machine learning methods have been effective in reconstructing various tracts with low false positive rates, reducing the systematic errors and biases that are prevalent with classical tractography methods, and improving generalizability to unseen data [56, 18]. Several reasons have been put forward to explain the success of machine learning methods. First, they avoid making rigid, ad-hoc, and possibly sub-optimal modeling assumptions or streamline propagation and stopping rules that are common in standard methods [56, 18, 238]. Some of the proposed machine learning methods use the dMRI data (rather than diffusion tensor or FOD estimates) as input, thereby sidestepping the unavoidable errors in fiber orientation estimation [56, 235, 59]. Moreover, they have the potential to learn tissue probabilities from dMRI data, rather than relying on the information provided by a registered anatomical MRI or ambiguous parameters such as FA. Machine learning methods can also learn the noise and artifacts from data, instead of imposing simplified noise models and assuming artifact-free scans [56]. Furthermore, these methods can incorporate estimation uncertainty into streamline tractography in a systematic way [238]. Machine learning techniques are also by far more flexible in incorporating non-local information, such as anatomical context, which can greatly improve tractography results [18]. An important piece of information that can be easily incorporated into most machine learning models is the direction of preceding streamline tracing steps. The direction of the last step indicates the streamline slope and the direction of the last two steps indicate its curvature. This information has been useful in improving tractography accuracy [234]. Incorporation of the direction of preceding steps can also improve the tractography in voxels with crossing fibers and bottleneck regions [237, 234].
@@ IV-F4 Tractography post-processing
Machine learning methods have also been used to enhance the generated tractograms via post-processing operations such as identification of false positives or anatomically implausible streamlines. Deep learning has shown great success at this task as well. An example of a supervised method is the work of Astolfi [252], where an existing method [253] is used to generate labels on training data based on anatomical priors. A limitation of supervised approaches is that it is hard to obtain training data that can adequately represent the complete range of invalid streamlines [254]. Unsupervised methods sidestep this limitation. An example of unsupervised techniques is the auto-encoder method proposed by Legarreta [255]. This method uses unlabeled whole-brain tractograms to learn a low-dimensional representation of valid streamlines, which can then be used to detect spurious or anatomically invalid streamlines in the test data. Both supervised and unsupervised approaches have reported remarkable success.
@@ IV-J Segmentation
More recently, Golkov et al. [16] extended their deep learning tissue microstructure estimation technique to develop a segmentation method that mapped the dMRI signal in each voxel to a probability for white matter (WM), gray mater (GM), cerebrospinal fluid (CSF), and MS lesion. Their simple and flexible approach allowed them to include an extra FLAIR channel, which could be especially useful for MS lesion segmentation. Their model accurately segmented the brain tissue, and detected MS lesions with area under the receiver operator characteristic curve (AUC) in the range 0.88-0.94. The same authors also proposed a variational auto-encoder model for novelty detection based on q-space dMRI data and showed that it could accurately segment MS lesions with an AUC of approximately 0.90 [350]. Another study used DTI and DKI features as input to a multi-view CNN to achieve accurate segmentation of WM, GM, and CSF [351]. The model was trained on HCP-quality data, where the anatomical and dMRI data could be registered precisely. The trained model achieved high segmentation accuracy on clinical dMRI acquisitions. The same research team trained a deep learning model to parcellate the brain into 101 regions using DTI-derived parameter maps as input [352]. The authors showed the utility of the new method in improving tract-specific streamline bundle extraction from whole-brain tractograms. A recent study addressed the challenging task of fetal brain segmentation in dMRI [353]. This is a particularly difficult application due to the low contrast of fetal brain dMRI. The authors used DTI-derived parameters as the input to their model, which was based on vision transformers. They proposed a novel self-supervised learning approach that enabled them to train the model on a large unlabeled dataset and a smaller labeled dataset. This method achieved a mean DSC of 0.84-0.90 on different tissue types. Experiments showed that this method drastically improved fetal brain tractography. For isointense stage of brain development where white matter and gray matter have similar intensity in T1- and T2-weighted images (approximately between 6 and 8 months), one study used an additional FA channel to accurately segment the brain tissue with an FCN [354]. Segmentation of the intracranial volume also know as brain masking or skull stripping, which is typically performed as a pre-processing step, has also been addressed with machine learning methods [355, 356, 357, 358].
@@ VI Conclusions
This research was supported in part by the National Institute of Neurological Disorders and Stroke and Eunice Kennedy Shriver National Institute of Child Health and Human Development of the National Institutes of Health (NIH) under award numbers R01HD110772 and R01NS128281. The content of this publication is solely the responsibility of the author and does not necessarily represent the official views of the NIH.
Declaration of Competing Interests
The author has no conflicts of interest to declare.

### v3

@@ Abstract
Diffusion-weighted magnetic resonance imaging (dMRI) of the brain offers unique capabilities including noninvasive probing of tissue microstructure and structural connectivity. It is widely used for clinical assessment of disease and injury, and for neuroscience research. Analyzing the dMRI data to extract useful information for medical and scientific purposes can be challenging. The dMRI measurements may suffer from strong noise and artifacts, and may exhibit high inter-session and inter-scanner variability in the data, as well as inter-subject heterogeneity in brain structure. Moreover, the relationship between measurements and the phenomena of interest can be highly complex. Recent years have witnessed increasing use of machine learning methods for dMRI analysis. This manuscript aims to assess these efforts, with a focus on methods that have addressed data preprocessing and harmonization, microstructure mapping, tractography, and white matter tract analysis. We study the main findings, strengths, and weaknesses of the existing methods and suggest topics for future research. We find that machine learning may be exceptionally suited to tackle some of the difficult tasks in dMRI analysis. However, for this to happen, several shortcomings of existing methods and critical unresolved issues need to be addressed. There is a pressing need to improve evaluation practices, to increase the availability of rich training datasets and validation benchmarks, as well as model generalizability, reliability, and explainability concerns.
@@ I-A Background and motivation
Classical dMRI analysis methods have appropriately been based on conventional signal processing, biophysical modeling, and numerical optimization techniques. Meanwhile, many studies have advocated for machine learning and data-driven techniques [16, 17, 18] motivated by the opportunities to exploit advances in hardware, new software libraries, and models learned from data. These techniques have become more popular in recent years. This trend has been driven by more powerful machine learning models (mostly based on deep neural networks) and a greater appreciation of the power and flexibility of these methods. Overall, these methods have been shown to possess the potential to improve the speed, accuracy, and reproducibility of various computations in dMRI analysis such as data pre-processing [19] and harmonization [20], tissue microstructure mapping [16, 21, 22], tractography [18], tract-specific analysis [17], and population studies [23]. It appears that recent studies herald a new generation of dMRI analysis methods that may significantly complement, if not at least in some cases entirely replace, the more conventional techniques. Therefore, it is good time to review, summarize, and critically assess the achievements of these works, highlight their shortcomings and limitations, and point out future work that may contribute to this new field in dMRI.
@@ I-B Scope and organization of this manuscript
The search to find the relevant works to be included in this paper was conducted in PubMed and Google Scholar. The initial search was performed in June 2023 and then repeated in June 2024. The search terms included “diffusion MRI” and “machine learning” or “deep learning”. More detailed searches were performed in reputable journals (e.g., Medical Image Analysis and IEEE Transactions on Medical Imaging) and conferences (e.g., MICCAI) that have a focus on this topic. Given the wide scope of work in this domain, we sought to focus on papers with interesting methodological advances or extensive experimental results.
This paper starts with a brief description of challenges in dMRI data analysis (Section II) and general overview of the reasons why machine learning may be effective in addressing these challenges (Section III). Section IV reviews prior works that have used machine learning in dMRI. The main applications considered include data preprocessing and quality enhancement, data harmonization, estimation of tissue microstructure and fiber orientation, tractography, tract-specific analysis, registration and segmentation. Figure 1 shows the outline of that section. Following that, Section V discusses the main technical considerations, challenges, and open questions. Some of the main topics discussed in that section include validation approaches, inherent limitations of machine learning, data, ground truth, and modeling considerations, as well as model explainability, uncertainty, and reliability concerns. Figure 2 shows the outline of that section. A short Conclusion section will present the closing remarks.
@@ II The challenging nature of dMRI analysis
The challenges and pitfalls of analyzing the dMRI data have been discussed in dedicated publications [9, 15, 24]. A primary source of difficulty in dMRI analysis is measurement noise and artifacts [9, 25]. Measurement noise can significantly alter the analysis results even in relatively simple computations such as diffusion tensor estimation [26, 27]. It can be difficult to suppress the noise and artifacts due to their complex distributions. The echo planar imaging method that is used to acquire dMRI measurements can give rise to eddy-current- and susceptibility-induced distortions. Subject motion during data acquisition is another persistent challenge in MRI. These factors can significantly impact the accuracy of quantitative tissue microstructure mapping and structural connectivity analysis [28, 29, 30]. Furthermore, the relationship between the dMRI measurements and the underlying microstructure of neuronal tissue is complex [31, 32]. Another source of difficulty is the measurement requirements of advanced dMRI models. Estimation of multi-compartment models of tissue microstructure and of complex white matter fiber configurations requires densely-sampled q-space data. Models often seek to represent properties of brain tissue with parameters corresponding to aspects of microstructure. Classical signal processing approaches are concerned with characterizing the identifiability of model parameters, and with characterizing any potential bias or variance in estimated values of parameters. For some model parameters to be identifiable, imaging strategies may need to be carefully chosen to obtain suitable measurements. Obtaining such measurements may require long scan times, leading to a tradeoff between the acquisition duration and the phenomena captured by a model. Even with a reasonably large number of measurements, standard computational methods can produce erroneous results [33]. As only one example, conventional linear and non-linear least-squares techniques for diffusion kurtosis imaging (DKI [34]) tend to produce physically implausible results [35, 36]. Streamline tractography and tract-specific analysis face inherent challenges and ambiguities such as fiber crossings and bottlenecks [24, 37, 38, 39], which can give rise to high false positive rates and erroneous results. Cross-subject comparisons and population studies can also be significantly hampered by high inter-scanner data heterogeneity [40, 41]. Furthermore, most applications are also challenged by paucity of ground truth and lack of universally accepted performance metrics.
@@ IV-F Tractography and structural connectivity
Machine learning may offer a framework for developing better tractography algorithms. They are model-free and allow for a seamless integration of anatomical priors and other sources of information that may be useful for tractography. Some of these information, such as the direction of previous tractography steps, have been used in conventional tractography methods in the past [227, 228]. Nonetheless, machine learning methods offer higher flexibility in integrating various inputs in a unified model and to optimize the model to reduce the tractography error with respect to all inputs jointly. As a result, they have the potential to have lower false positive rates, to be more robust to measurement noise, and to overcome the inherent biases of classical tractography algorithms [18, 229, 230]. Although various methods such as self-organizing maps [231] and random forests [229] have been successfully used for tractography, machine learning-based tractography techniques have increasingly relied on neural networks [18, 229, 232, 233]. Table VI lists some of the works that have developed machine learning methods for streamline tractography. We summarize the main technical aspects of these methods and their reported results below.
@@
Jörgens et al. [236] MLP (D) Raw diffusion signal & directions of the two preceding streamline steps. This work reports extensive experiments to determine the optimal input and network architectures.
Wegmayr et al. [59] MLP (D) Raw diffusion signal in a $3^{3}$-voxel neighborhood & directions of the two preceding streamline steps. Better results than standard tractography methods when trained on small datasets. The method is robust to noisy training data.
Cai et al. [61] Convolutional-recurrent neural network Anatomical MRI and streamline memory. This work has reported that the variability in tractography and structural connectivity metrics computed based on T1-weighted MRI is similar to that for standard dMRI-based results.
Liu et al. [237] Convolutional, attention, and MLP modules FOD, brain parcellation and tissue segmentation maps, prior fixel map & directions of six preceding steps. This method was designed for fetal brain tractography. Experiments showed superior ability to reconstruct various white matter tracts compared with existing methods.
Track-to-Learn [238] Deep reinforcement learning (D & P) FOD at the current voxel and its six neighbors & directions of the past four streamline steps. A reinforcement learning approach can be competitive with supervised learning methods while also being more generalizable to unseen or out-of-distribution data.
@@
Entrack [239] MLP (P) Directions of the three last streamline steps & FOD, represented as SH of order 4, in a voxel neighborhood of size 3 voxels. The method achieves competitive results compared with a range of conventional and machine learning methods on independent synthetic test data.
@@ IV-F3 Evaluations and main results
Machine learning methods have been effective in reconstructing various tracts with low false positive rates, reducing the systematic errors and biases that are prevalent with classical tractography methods, and improving generalizability to unseen data [56, 18]. Several reasons have been put forward to explain the success of machine learning methods. First, they avoid making rigid, ad-hoc, and possibly sub-optimal modeling assumptions or streamline propagation and stopping rules that are common in standard methods [56, 18, 239]. Some of the proposed machine learning methods use the dMRI data (rather than diffusion tensor or FOD estimates) as input, thereby sidestepping the unavoidable errors in fiber orientation estimation [56, 235, 59]. Moreover, they have the potential to learn tissue probabilities from dMRI data, rather than relying on the information provided by a registered anatomical MRI or ambiguous parameters such as FA. Machine learning methods can also learn the noise and artifacts from data, instead of imposing simplified noise models and assuming artifact-free scans [56]. Furthermore, these methods can incorporate estimation uncertainty into streamline tractography in a systematic way [239]. Machine learning techniques are also by far more flexible in incorporating non-local information, such as anatomical context, which can greatly improve tractography results [18]. An important piece of information that can be easily incorporated into most machine learning models is the directions of preceding streamline tracing steps. The direction of the last step indicates the streamline slope and the directions of the last two steps indicate its curvature. This information has been useful in improving tractography accuracy [234]. Incorporation of the direction of preceding steps can also improve the tractography in voxels with crossing fibers and bottleneck regions [238, 234].
@@ IV-F4 Tractography post-processing
Machine learning methods have also been used to enhance the generated tractograms via post-processing operations such as identification of false positives or anatomically implausible streamlines. Deep learning has shown great success in this task as well. An example of a supervised method is the work of Astolfi [253], where an existing method [254] is used to generate labels on training data based on anatomical priors. A limitation of supervised approaches is that it is hard to obtain training data that can adequately represent the complete range of invalid streamlines [255]. Unsupervised methods sidestep this limitation. An example of unsupervised techniques is the auto-encoder method proposed by Legarreta [256]. This method uses unlabeled whole-brain tractograms to learn a low-dimensional representation of valid streamlines, which can then be used to detect spurious or anatomically invalid streamlines in the test data. Both supervised and unsupervised approaches have reported remarkable success. A generative method based on autoencoders has recently been propsoed to improve the streamline density for tracts that are hard to trace with conventional propagation methods [257]. Using anatomically valid training streamlines, it learns a latent space representation that can be subsequently used as a “streamline yard” to synthesize new streamlines for sparsely-populated tracts in a test sample. The synthesized streamlines are accepted or rejected based on a set of criteria related to streamline geometry, connectivity, and agreement with local fiber orientation directions. Experiments with phantom and real data showed that this method significantly improved tract coverage.
@@ IV-J Segmentation
More recently, Golkov et al. [16] extended their deep learning tissue microstructure estimation technique to develop a segmentation method that mapped the dMRI signal in each voxel to a probability for white matter (WM), gray mater (GM), cerebrospinal fluid (CSF), and MS lesion. Their simple and flexible approach allowed them to include an extra FLAIR channel, which could be especially useful for MS lesion segmentation. Their model accurately segmented the brain tissue, and detected MS lesions with area under the receiver operator characteristic curve (AUC) in the range 0.88-0.94. The same authors also proposed a variational auto-encoder model for novelty detection based on q-space dMRI data and showed that it could accurately segment MS lesions with an AUC of approximately 0.90 [352]. Another study used DTI and DKI features as input to a multi-view CNN to achieve accurate segmentation of WM, GM, and CSF [353]. The model was trained on HCP-quality data, where the anatomical and dMRI data could be registered precisely. The trained model achieved high segmentation accuracy on clinical dMRI acquisitions. The same research team trained a deep learning model to parcellate the brain into 101 regions using DTI-derived parameter maps as input [354]. The authors showed the utility of the new method in improving tract-specific streamline bundle extraction from whole-brain tractograms.
A recent study addressed the challenging task of fetal brain segmentation in dMRI [355]. This is a particularly difficult application due to the low contrast of fetal brain dMRI. The authors used DTI-derived parameters as the input to their model, which was based on vision transformers. They proposed a novel self-supervised learning approach that enabled them to train the model on a large unlabeled dataset and a smaller labeled dataset. This method achieved a mean DSC of 0.84-0.90 on different tissue types. Experiments showed that this method drastically improved fetal brain tractography. Building upon their improved tractography results, the same research team has developed a multi-task learning method that segments the brain tissue, white matter tracts, and gray matter regions [356]. For isointense stage of brain development where white matter and gray matter have similar intensity in T1- and T2-weighted images (approximately between 6 and 8 months), one study used an additional FA channel to accurately segment the brain tissue with an FCN [357]. Segmentation of the intracranial volume also know as brain masking or skull stripping, which is typically performed as a pre-processing step, has also been addressed with machine learning methods [358, 359, 360, 361].
@@ VI Conclusions
This research was supported in part by the National Institute of Neurological Disorders and Stroke and Eunice Kennedy Shriver National Institute of Child Health and Human Development of the National Institutes of Health (NIH) under award numbers R01HD110772 and R01NS128281, and in part by NIH awards S10 OD025111 and R01 LM013608. The content of this publication is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.